OpenStudy (parissfrance09):

Please Help Will Medal :) An item is priced at $14.82. If the sales tax is 4%, what does the item cost including sales tax? A.$30.23 B.$0.59 C.$15.41 D.$20.75

OpenStudy (shamim):

Tax=14.82*4%=?

OpenStudy (parissfrance09):

0.5928??

OpenStudy (shamim):

Right!!!

OpenStudy (parissfrance09):

so the answer is B??

OpenStudy (shamim):

Total cost=14.82+0.5928=?

OpenStudy (parissfrance09):

15.4128

OpenStudy (shamim):

Right!!!

OpenStudy (parissfrance09):

so my answer is c??

OpenStudy (shamim):

Ya c

OpenStudy (parissfrance09):

thank you

OpenStudy (shamim):

U r most welcome!!!!!!!